OpenStudy (anonymous):

What volume of a 3.5 M HCl is required to completely neutralize 50.0 ml of a 2.0 M NaOH?

OpenStudy (somy):

by M u mean concentration?

OpenStudy (somy):

if so then first u need to write balanced equation NaOH + HCl -> NaCl + H2O for NaOH we know volume and concentration so we use formula concentration = mole/ volume we need to find mole, thus we make it as subject mole = concentration*volume mole = 2.0*50 = 100moles of NaOH reacted with HCl from equation you can see that the ratio between NaOH and HCl is 1:1 thus mole of HCl is also 100 mole so for HCl we already know concentration and now mole also we again use concentration = mole/volume formula this time we'll make volume as subject so volume = mole/concentration volume = 100 / 3.5 = 28.57 ml i hope u get the working and next time you can do it on your own :)